# G4Splat: Geometry-Guided Gaussian Splatting with Generative Prior

### v1

@@ Abstract
Despite recent advances in leveraging generative prior from pre-trained diffusion models for 3D scene reconstruction, existing methods still face two critical limitations. First, due to the lack of reliable geometric supervision, they struggle to produce high-quality reconstructions even in observed regions, let alone in unobserved areas. Second, they lack effective mechanisms to mitigate multi-view inconsistencies in the generated images, leading to severe shape–appearance ambiguities and degraded scene geometry. In this paper, we identify accurate geometry as the fundamental prerequisite for effectively exploiting generative models to enhance 3D scene reconstruction. We first propose to leverage the prevalence of planar structures to derive accurate metric-scale depth maps, providing reliable supervision in both observed and unobserved regions. Furthermore, we incorporate this geometry guidance throughout the generative pipeline to improve visibility mask estimation, guide novel view selection, and enhance multi-view consistency when inpainting with video diffusion models, resulting in accurate and consistent scene completion. Extensive experiments on Replica, ScanNet++, and DeepBlending show that our method consistently outperforms existing baselines in both geometry and appearance reconstruction, particularly for unobserved regions. Moreover, our method naturally supports single-view inputs and unposed videos, with strong generalizability in both indoor and outdoor scenarios with practical real-world applicability. The project page is available at https://dali-jack.github.io/g4splat-web/.
Figure 1: We propose G4Splat, which integrates accurate geometry guidance with generative prior to enhance 3D scene reconstruction. Our method significantly improves geometry and appearance reconstruction quality in both observed and unobserved regions (black areas in baselines), and generalizes well across diverse scenarios, including single-view scenes and unposed videos.
@@ 1 Introduction
In this paper, we argue that robust geometry guidance is fundamental to effectively employing the power of generative models and realizing high-quality reconstruction. To this end, we introduce G4Splat, which first leverages the prevalence of planar structures in man-made environments, consistent with the Manhattan world assumption (Coughlan & Yuille, 1999), to derive scale-accurate geometric constraints. Unlike matching-based methods (Wang et al., 2024a; Guédon et al., 2025) that often fail in non-overlapping regions, planar surfaces allow depth extrapolation: a 3D plane can be reliably estimated from partial depth observations and then extended across the entire surface. We leverage this property by aligning global 3D planes to obtain scale-accurate depth in planar regions and linearly adjusting monocular depth for non-planar regions, resulting in plane-aware depth maps that provide accurate geometric supervision in both observed and unobserved regions.
@@
Experiments on Replica (Straub et al., 2019), ScanNet++ (Yeshwanth et al., 2023), and DeepBlending (Hedman et al., 2018) demonstrate that our method consistently outperforms all baselines in both geometric and appearance reconstruction, with particularly strong improvements in unobserved regions. Ablation studies further confirm the effectiveness of each proposed component. Moreover, our approach supports reconstruction from casually captured videos or text-generated single images, as illustrated in Figs.˜1 and 5, with significant potential in practical downstream applications in embodied AI (Huang et al., 2024b; 2025b; Jia et al., 2024), robotics (Li et al., 2024c; e; Zhi et al., 2025), and more (Jiang et al., 2024; Wang et al., 2024b).
@@ 2.1 Sparse-View 3DGS
3D Gaussian Splatting (3DGS)(Kerbl et al., 2023; Lu et al., 2024a; Yu et al., 2024c) has demonstrated impressive reconstruction quality and training efficiency, but its performance deteriorates significantly when trained with only a few input views. To address this limitation, numerous approaches(Kumar & Vats, 2025; Zhang et al., 2024b; Wu et al., 2025b; Hong et al., 2025; Huang et al., 2025a; Xu et al., 2024; Paliwal et al., 2024; Zhao et al., 2024; 2025a; Bao et al., 2024; Liu et al., 2024f; Yin et al., 2024; Xiao et al., 2024; Lu et al., 2024b; Zheng et al., 2025) have been proposed. For example, DNGaussian (Li et al., 2024a) and FSGS (Zhu et al., 2024) incorporate depth regularization to suppress floaters in visible regions; yet, the scale ambiguity inherent in monocular depth estimators (Yang et al., 2024) prevents them from providing reliable geometric supervision. MAtCha (Guédon et al., 2025) attempts to overcome this limitation by introducing scale-accurate depth from structure-from-motion (SfM) methods (Wang et al., 2024a; Duisterhof et al., 2025). Nevertheless, it still struggles to reconstruct non-overlapping regions between input views. To address this issue, we propose to leverage the inherent extensibility of plane representations that propagate accurate depth estimates from overlapping regions to non-overlapping or even unobserved regions.
@@ 2.2 Generative Prior for 3DGS
Recent studies (Poole et al., 2022; Xiong et al., 2023; Weber et al., 2024; Wu et al., 2024; Liu et al., 2024e; Melas-Kyriazi et al., 2024; Shih et al., 2024; Paul et al., 2024; Cai et al., 2024; Chen et al., 2024b; Yu et al., 2024a; Ni et al., 2025) have demonstrated the effectiveness of leveraging diffusion models (Rombach et al., 2022; Liu et al., 2023) to provide powerful priors for 3D reconstruction. More recent approaches (Liu et al., 2024a; b; d; Zhao et al., 2025b; Gao et al., 2024; Zhou et al., 2025; Bao et al., 2025; Wu et al., 2025a; c; Fischer et al., 2025; Yin et al., 2025; Zhong et al., 2025) further employ video diffusion models (Blattmann et al., 2023; Yu et al., 2024b) to enhance cross-view consistency. However, since these models primarily rely on initial reconstructions to apply their generative power, they still suffer from numerous floating Gaussian artifacts and low-quality 3D geometry in their results, particularly in the inpainted regions, due to insufficient geometric constraints under sparse observations. To address this issue, we first apply scale-accurate depth supervision for both observed and non-observed regions to provide a robust geometric foundation, and integrate geometric guidance throughout the video diffusion models, leading to significantly improved reconstruction of both geometry and appearance.
@@ 2.3 Plane Assumption in Reconstruction
The Manhattan-world assumption (Coughlan & Yuille, 1999), and in particular the plane assumption, has been widely adopted in the reconstruction of man-made environments. Several methods (Liu et al., 2024c; Guo et al., 2024; Mazur et al., 2024; Liu et al., 2025b; Pataki et al., 2025) leverage this assumption in SfM and SLAM to improve matching accuracy, while other plane reconstruction approaches (Liu et al., 2019; Agarwala et al., 2022; Xie et al., 2022; Tan et al., 2023; Watson et al., 2024; Ye et al., 2025; Liu et al., 2025a) directly fit a set of planes to model indoor scenes. More recent studies (Guo et al., 2022; Li et al., 2024b; Chen et al., 2024a; Shi et al., 2025) incorporate the plane assumption to optimize 3D neural implicit representations (Mildenhall et al., 2020; Kerbl et al., 2023). For instance, GeoGaussian (Li et al., 2024d) and IndoorGS (Ruan et al., 2025) impose local planar constraints to regulate Gaussian splitting and movement, whereas PlanarSplatting (Tan et al., 2025) reconstructs 3D scenes as planar primitives directly from multi-view images. Different from these methods, our approach leverages the plane assumption to extract scale-accurate depth, which not only provides geometric guidance for optimizing the Gaussian representation but also facilitates the integration of generative prior, ultimately enabling precise scene reconstruction for both planar and non-planar structures across observed and unobserved regions.
@@ 3.2 Plane-Aware Geometry Modeling
Inspired by prior work (Mazur et al., 2024; Ye et al., 2025), we assume that planar regions within an image exhibit consistent normal directions, smooth geometry, and similar semantics. Based on this assumption, we extract plane masks from images by combining normal maps with instance masks generated by SAM (Kirillov et al., 2023). Specifically, we apply K-means clustering to the normal map to obtain regions with coherent surface orientations. These regions are subsequently filtered using SAM masks, and only those assigned the same instance label and exceeding a predefined size threshold are treated as valid 2D plane masks. Example results are shown in Fig.˜3(a).
@@
where $\mathbf{n}_{k}\in\mathbb{R}^{3}$ is the unit normal vector and $d_{k}\in\mathbb{R}$ is the offset. To robustly estimate the plane parameters, we select a subset of high-confidence points $\mathcal{P}_{k}^{\text{conf}}\subset\mathcal{P}_{k}$, defined as those observed in at least two different views. A 3D plane is then fitted to $\mathcal{P}_{k}^{\text{conf}}$ by RANSAC (Fischler & Bolles, 1981):
@@
For non-planar regions in $I^{v}$, we adopt a pre-trained monocular depth estimator (Yang et al., 2024) to predict a relative depth map $\hat{D}^{v}$. This relative map is aligned to an absolute scale using the already computed plane region depths $\{D_{i}^{v}\}$ via a linear transformation:
@@
where the scale $a_{v}$ and offset $b_{v}$ are estimated by least-squares fitting over pixels belonging to planar regions. The resulting depth map $D^{v}$ integrates geometry-consistent plane depths with refined monocular predictions in non-planar areas, yielding a complete and scale-accurate plane-aware depth representation for each view. As illustrated in Fig.˜3(a), this plane-aware depth significantly alleviates errors in non-overlapping regions when compared to MAtCha (Guédon et al., 2025).
@@ 3.3 Geometry-guided Generative Pipeline
Despite this joint inference, the inpainted results $\{\hat{I}^{v}\}$ from existing generative models still exhibit multi-view inconsistencies, which can introduce artifacts during training (Zhong et al., 2025). To mitigate this, for each global 3D plane, we primarily rely on the color supervision from the view that observes the plane most completely, thereby reducing cross-view conflicts. A more detailed discussion is provided in Section˜C.3.
Figure 4: Qualitative comparison from 5 input views. Our approach achieves better appearance and geometry reconstruction with fewer Gaussian floaters in both observed and unobserved regions.
@@ 3.4 Overall Training Strategy
In the initialization stage, we first apply chart alignment in MAtCha to obtain an initial depth map for each input view. From these depth maps, we estimate global 3D planes and compute plane-aware depth maps, as described in Section˜3.2. The Gaussian parameters are then initialized from the point cloud of the depth maps and trained with plane-aware depth maps, producing a baseline model with accurate geometry in the regions observed by the input views.
@@ 4 Experiments
We evaluate G4Splat on both geometric and appearance reconstruction in sparse-view 3D scenarios. Additionally, we present more experimental results in Appendix˜A, failure cases and discuss the method’s limitations in Appendix˜D.
@@ 4.1 Settings
Datasets and Metrics
We evaluate our method on both synthetic and real-world datasets. The synthetic dataset is Replica (Straub et al., 2019), which contains 8 indoor scenes. The real-world datasets include 6 scenes from ScanNet++(Yeshwanth et al., 2023) and 3 scenes from DeepBlending(Hedman et al., 2018). For quantitative comparisons, we evaluate Chamfer Distance (CD), F-Score, and Normal Consistency (NC) as reconstruction metrics, and PSNR, SSIM, and LPIPS as rendering metrics. More details are provided in Sections˜C.5 and C.6.
@@
We compare our method with several representative baselines. These include the classical Gaussian Splatting approaches 3DGS (Kerbl et al., 2023) and 2DGS (Huang et al., 2024a), as well as state-of-the-art sparse-view 3DGS methods FSGS (Zhu et al., 2024), InstantSplat (Fan et al., 2024), and MAtCha (Guédon et al., 2025). We further evaluate against recent approaches that integrate diffusion models into sparse-view 3DGS reconstruction, namely GenFusion (Wu et al., 2025c), Difix3D+ (Wu et al., 2025a) and GuidedVD (Zhong et al., 2025). In addition, we implement a variant of 2DGS augmented with the See3D (Ma et al., 2025). Notably, all baselines are augmented with MASt3R-SfM (Duisterhof et al., 2025) to improve robustness in sparse-view scenarios.

### v2

@@ Abstract
Despite recent advances in leveraging generative prior from pre-trained diffusion models for 3D scene reconstruction, existing methods still face two critical limitations. First, due to the lack of reliable geometric supervision, they struggle to produce high-quality reconstructions even in observed regions, let alone in unobserved areas. Second, they lack effective mechanisms to mitigate multi-view inconsistencies in the generated images, leading to severe shape–appearance ambiguities and degraded scene geometry. In this paper, we identify accurate geometry as the fundamental prerequisite for effectively exploiting generative models to enhance 3D scene reconstruction. We first propose to leverage the prevalence of planar structures to derive accurate metric-scale depth maps, providing reliable supervision in both observed and unobserved regions. Furthermore, we incorporate this geometry guidance throughout the generative pipeline to improve visibility mask estimation, guide novel view selection, and enhance multi-view consistency when inpainting with video diffusion models, resulting in accurate and consistent scene completion. Extensive experiments on Replica, ScanNet++, DeepBlending and Mip-NeRF 360 show that our method consistently outperforms existing baselines in both geometry and appearance reconstruction, particularly for unobserved regions. Moreover, our method naturally supports single-view inputs and unposed videos, with strong generalizability in both indoor and outdoor scenarios with practical real-world applicability. Project page: https://dali-jack.github.io/g4splat-web/.
Figure 1: We propose G4Splat, which integrates accurate geometry guidance with generative prior to enhance 3D scene reconstruction. Our method significantly improves geometry and appearance reconstruction quality in both observed and unobserved regions (black areas in baselines), and generalizes well across diverse scenarios, including single-view images and unposed videos.
@@ 1 Introduction
In this paper, we argue that robust geometry guidance is fundamental to effectively employing the power of generative models and realizing high-quality reconstruction. To this end, we introduce G4Splat, which first leverages the prevalence of planar structures in man-made environments, consistent with the Manhattan world assumption (Coughlan and Yuille, 1999), to derive scale-accurate geometric constraints. Unlike matching-based methods (Wang et al., 2024a; Guédon et al., 2025) that often fail in non-overlapping regions, planar surfaces allow depth extrapolation: a 3D plane can be reliably estimated from partial depth observations and then extended across the entire surface. We leverage this property by aligning global 3D planes to obtain scale-accurate depth in planar regions and linearly adjusting monocular depth for unobserved non-planar regions, resulting in plane-aware depth maps that provide accurate geometric supervision in both observed and unobserved regions.
@@
Experiments on Replica (Straub et al., 2019), ScanNet++ (Yeshwanth et al., 2023), DeepBlending (Hedman et al., 2018) and Mip-NeRF 360 (Barron et al., 2022) demonstrate that our method consistently outperforms all baselines in both geometric and appearance reconstruction, with particularly strong improvements in unobserved regions. Ablation studies further confirm the effectiveness of each proposed component. Moreover, our approach supports reconstruction from casually captured videos or text-generated single images, as illustrated in Figs.˜1 and 5, with significant potential in practical downstream applications in embodied AI (Huang et al., 2024b; 2025c; Jia et al., 2024), robotics (Li et al., 2024c; e; Zhi et al., 2025), and more (Jiang et al., 2024; Wang et al., 2024b).
@@ 2.1 Sparse-View 3DGS
3D Gaussian Splatting (3DGS)(Kerbl et al., 2023; Lu et al., 2024a; Yu et al., 2024c) has demonstrated impressive reconstruction quality and training efficiency, but its performance deteriorates significantly when trained with only a few input views. To address this limitation, numerous approaches(Kumar and Vats, 2025; Zhang et al., 2024b; Wu et al., 2025b; Hong et al., 2025; Huang et al., 2025b; Xu et al., 2024; Paliwal et al., 2024; Zhao et al., 2024; 2025a; Bao et al., 2024; Liu et al., 2024f; Yin et al., 2024; Xiao et al., 2024; Lu et al., 2024b; Zheng et al., 2025) have been proposed. For example, DNGaussian (Li et al., 2024a) and FSGS (Zhu et al., 2024) incorporate depth regularization to suppress floaters in visible regions; yet, the scale ambiguity inherent in monocular depth estimators (Yang et al., 2024) prevents them from providing reliable geometric supervision. MAtCha (Guédon et al., 2025) attempts to overcome this limitation by introducing scale-accurate depth from structure-from-motion (SfM) methods (Wang et al., 2024a; Duisterhof et al., 2025). Nevertheless, it still struggles to reconstruct non-overlapping regions between input views. To address this issue, we propose to leverage the inherent extensibility of plane representations that propagate accurate depth estimates from overlapping regions to non-overlapping or even unobserved regions.
@@ 2.2 Generative Prior for 3DGS
Recent studies (Poole et al., 2022; Xiong et al., 2023; Weber et al., 2024; Wu et al., 2024; Liu et al., 2024e; Melas-Kyriazi et al., 2024; Shih et al., 2024; Paul et al., 2024; Cai et al., 2024; Chen et al., 2024b; Yu et al., 2024a; Ni et al., 2025) have demonstrated the effectiveness of leveraging diffusion models (Rombach et al., 2022; Liu et al., 2023) to provide powerful priors for 3D reconstruction. Some methods (Raj et al., 2025; Huang et al., 2025a; Chang et al., 2025) attempt to introduce 3D geometry priors into the reconstruction process. More recent approaches (Liu et al., 2024a; b; d; Zhao et al., 2025b; Gao et al., 2024; Zhou et al., 2025; Bao et al., 2025; Wu et al., 2025a; c; Fischer et al., 2025; Yin et al., 2025; Zhong et al., 2025) further employ video diffusion models (Blattmann et al., 2023; Yu et al., 2024b) to enhance cross-view consistency. However, since these models primarily rely on initial reconstructions to apply their generative power, they still suffer from numerous floating Gaussian artifacts and low-quality 3D geometry in their results, particularly in the inpainted regions, due to insufficient geometric constraints under sparse observations. To address this issue, we first apply scale-accurate depth supervision for both observed and non-observed regions to provide a robust geometric foundation, and integrate geometric guidance throughout the video diffusion models, leading to significantly improved reconstruction of both geometry and appearance.
@@ 2.3 Plane Assumption in Reconstruction
The Manhattan-world assumption (Coughlan and Yuille, 1999), and in particular the plane assumption, has been widely adopted in the reconstruction of man-made environments. Several methods (Liu et al., 2024c; Guo et al., 2024; Mazur et al., 2024; Liu et al., 2025b; Pataki et al., 2025) leverage this assumption in SfM and SLAM to improve matching accuracy, while other plane reconstruction approaches (Liu et al., 2019; Agarwala et al., 2022; Xie et al., 2022; Tan et al., 2023; Watson et al., 2024; Ye et al., 2025; Liu et al., 2025a) directly fit a set of planes to model indoor scenes. More recent studies (Guo et al., 2022; Li et al., 2024b; Chen et al., 2024a; Shi et al., 2025) incorporate the plane assumption to optimize 3D neural implicit representations (Mildenhall et al., 2020; Kerbl et al., 2023). For instance, GeoGaussian (Li et al., 2024d) and IndoorGS (Ruan et al., 2025) impose local planar constraints to regulate Gaussian splitting and movement, whereas PlanarSplatting (Tan et al., 2025) reconstructs 3D scenes as planar primitives directly from multi-view images. Different from these methods, our approach leverages the plane assumption to extract scale-accurate depth, which not only provides geometric guidance for optimizing the Gaussian representation but also facilitates the integration of generative prior, ultimately enabling precise scene reconstruction for both planar and non-planar structures across observed and unobserved regions.
@@ 3 Method
We propose G4Splat, a method that integrates accurate geometry guidance with generative priors to enhance 3D scene reconstruction. We begin by introducing the base model MAtCha (Guédon et al., 2025) and the overall training objective in Section˜3.1. Next, we present our plane-aware geometry modeling in Section˜3.2, followed by the geometry-guided generative pipeline in Section˜3.3. Finally, we describe the overall training strategy in Section˜3.4.
@@ 3.2 Plane-Aware Geometry Modeling
Inspired by prior work (Mazur et al., 2024; Ye et al., 2025), we assume that planar regions within an image exhibit consistent normal directions, smooth geometry, and similar semantics. Based on this assumption, we extract plane masks from images by combining instance masks from SAM (Kirillov et al., 2023) with normal maps generated from either depth-map gradients or a monocular predictor (Ye et al., 2024). Specifically, we apply K-means clustering to the normal map to obtain regions with coherent surface orientations. These regions are subsequently filtered using SAM masks, and only those assigned the same instance label and exceeding a predefined size threshold are treated as valid 2D plane masks. Example results are shown in Fig.˜3(a).
@@
where $\mathbf{n}_{k}\in\mathbb{R}^{3}$ is the unit normal vector and $d_{k}\in\mathbb{R}$ is the offset. To robustly estimate the plane parameters, we select a subset of high-confidence points $\mathcal{P}_{k}^{\text{conf}}\subset\mathcal{P}_{k}$, defined as those observed in at least two different views. A 3D plane is then fitted to $\mathcal{P}_{k}^{\text{conf}}$ by RANSAC (Fischler and Bolles, 1981):
@@
For the non-planar regions of $I^{v}$ that are visible in the input view, we retain the depth values estimated by MAtCha (Guédon et al., 2025); for those that are not visible, we adopt a pre-trained monocular depth estimator (Yang et al., 2024) to predict a relative depth map $\hat{D}^{v}$. This relative map is aligned to an absolute scale using the already computed plane region depths $\{D_{i}^{v}\}$ via a linear transformation:
@@
where the scale $a_{v}$ and offset $b_{v}$ are estimated by least-squares fitting over pixels belonging to planar regions. The resulting depth map $D^{v}$ integrates geometry-consistent plane depths with refined monocular predictions in unobserved non-planar areas, yielding a complete and scale-accurate plane-aware depth representation for each view. As illustrated in Fig.˜3(a), this plane-aware depth significantly alleviates errors in non-overlapping regions when compared to MAtCha.
@@ 3.3 Geometry-guided Generative Pipeline
Despite this joint inference, the inpainted results $\{\hat{I}^{v}\}$ from existing generative models still exhibit multi-view inconsistencies, which can introduce artifacts during training (Zhong et al., 2025). To address this issue, we supervise each region by primarily relying on color information from a single view. For planar regions, we choose the view that provides the most complete observation of the corresponding plane, ensuring consistent inpainting. For non-planar regions, we use the first view in which the region becomes visible, which helps reduce cross-view inconsistencies. A more detailed discussion is provided in Section˜C.3.
Figure 4: Qualitative comparison. Our approach achieves better appearance and geometry reconstruction with fewer Gaussian floaters in both observed and unobserved regions.
@@ 3.4 Overall Training Strategy
In the initialization stage, we first apply chart alignment in MAtCha to obtain an initial depth map for each input view. From these depth maps, we estimate global 3D planes and compute plane-aware depth maps, as described in Section˜3.2. The Gaussian parameters are then initialized from the resulting point cloud and optimized using these plane-aware depth maps, producing a baseline model with accurate geometry in the regions observed by the input views.
@@ 4 Experiments
We evaluate G4Splat on both geometric and appearance reconstruction in sparse-view 3D scenarios. Additionally, we present more experimental results in Appendix˜A, failure cases and discuss the method’s limitations in Appendix˜D. For more extensive qualitative results and video demonstrations, please refer to our project page.
@@ 4.1 Settings
Datasets
We evaluate our method on both synthetic and real-world datasets. The synthetic dataset is Replica (Straub et al., 2019), which contains 8 indoor scenes. The real-world datasets include 6 scenes from ScanNet++ (Yeshwanth et al., 2023), 3 scenes from DeepBlending (Hedman et al., 2018) and 9 scenes from Mip-NeRF 360 (Barron et al., 2022).
For ScanNet++, DeepBlending, and Replica, we uniformly sample 100 images per scene. For the former two, we randomly select 5 images as input views and use the remaining 95 for testing. For Replica, we conduct experiments with 5, 10, and 15 input views to assess performance under varying view sparsity (Section˜A.3), while maintaining a consistent set of 85 test views across all settings. For Mip-NeRF 360, we follow the 9-input-view configuration as established in prior work (Wu et al., 2024; Gao et al., 2024).
Metrics
Reconstruction quality is quantified using Chamfer Distance (CD), F-Score, and Normal Consistency (NC), while rendering performance is measured via PSNR, SSIM, and LPIPS. As the Mip-NeRF 360 dataset lacks ground-truth meshes, we evaluate only the rendering performance for those scenes; for the remaining three datasets, we report the full suite of metrics. Further details are provided in Sections˜C.5 and C.6.
@@
We compare our method with several representative baselines. These include the classical Gaussian Splatting approaches 3DGS (Kerbl et al., 2023) and 2DGS (Huang et al., 2024a), as well as state-of-the-art sparse-view 3DGS methods FSGS (Zhu et al., 2024), InstantSplat (Fan et al., 2024), and MAtCha (Guédon et al., 2025). We further evaluate against highly competitive recent approaches that integrate generative diffusion models into sparse-view 3DGS reconstruction, namely GenFusion (Wu et al., 2025c), Difix3D+ (Wu et al., 2025a) and GuidedVD (Zhong et al., 2025). In addition, we implement a variant of 2DGS augmented with the See3D (Ma et al., 2025). Notably, all baselines are augmented with MASt3R-SfM (Duisterhof et al., 2025) to provide robust geometric initialization and improve performance in sparse-view scenarios.
Table 2: Quantitative comparison from 9 input views on Mip-NeRF 360.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
MAtCha 16.44 0.453 0.389
GenFusion 17.82 0.452 0.439
Difix3D+ 16.28 0.389 0.426
GuidedVD 16.78 0.438 0.499
See3D 16.92 0.443 0.451
Ours 18.66 0.515 0.371